OpenStudy (anonymous):

The linear approximation for the function f(x)=x^5 near x=2 is 32+80(x-2) Check this. need help with B. A.) What number c will give the best quadratic approximation x^5 approximate to 32+80(x-2)+c(x-2)^2. With the help of myininaya I got c=80. B.)If this approximation is used for various x's in the interval [2,2.1], can you be certian that the error is no bigger than .05? Explain using Taylor's inequality (the Error Bound). C.) Graph equation Part A now that we have c. On interval [2,2.1]. note: I don't need someone to graph this, I just want to make sure I have right approach. Would i just pick an x on the interval and set the equation to y and solve? for a point

OpenStudy (anonymous):

What I had help with earlier was part A and we ended up taking derivative up to the 2nd degree. and f"(x) was 20x^3 or 160. the taylor poly centered at x=a so f"(a)/2 was 160/2 or 80=c.

zepdrix (zepdrix):

Were you able to understand the part before part A? The linear approximation for the function f(x)=x^5 near x=2 is 32+80(x-2) `Check this.`

OpenStudy (anonymous):

x^5=32+80(x-2) plug in f(2) 32=32+80(0) 32+32

OpenStudy (anonymous):

32=32*

zepdrix (zepdrix):

Hmm ok. And part A makes sense maybe? :x Hmm I'm not sure about this Error Bound nonsense XD lol lemme think..

OpenStudy (anonymous):

I'm looking at the formula and explanation in my book. should i type it for you?

zepdrix (zepdrix):

Sure :D

OpenStudy (anonymous):

note: [ ] will represent absolute value. and *n* is subscript assume that f^(n+1) (x) exists and is continuous. Let K be a number such that: [f(x)-T*n* (x)] is less than/equal to (K) (([x-a]^(n+1))/(n+1)!) . where T*n* (x) is the nth Taylor polynomial centered at x+a.

OpenStudy (anonymous):

*n* is subscript n

OpenStudy (anonymous):

absolute value also on the x-a in the numerator.

zepdrix (zepdrix):

Oh ok :x\[\Large \left|f(x)-T_n(x)\right|\quad\le\quad k\frac{|x-a|^{n+1}}{(n+1)!}\]

OpenStudy (anonymous):

nice. I have to find out how to do that ha

OpenStudy (anonymous):

nice an clean that way

OpenStudy (anonymous):

part a was a second degree so i guess it would be T*2*

OpenStudy (anonymous):

f(x) =2 n=2?

zepdrix (zepdrix):

Grr .. brain.. esploding :c

zepdrix (zepdrix):

Hmm I don't think I can make sense of this :( At least not in a reasonable amount of time, lol. I found some good videos on the topic though D: Maybe they'll help. http://www.youtube.com/watch?v=wgkRH5Uoavk http://www.youtube.com/watch?v=yUUPP70Fhpo

zepdrix (zepdrix):

Imma run to the grocery store real quick :D

OpenStudy (anonymous):

okay. thank you for the help.